ind the general solution of the given system.
dx /dt = 3x − y − z
dy /dt = x + y − z
dz /dt = x − y + z x(t), y(t), z(t) =
1
Expert's answer
2020-11-19T16:44:24-0500
"Solution"

In matrix from, the system is


"\frac{d}{dt}\begin{bmatrix}x \\ y \\ z \end{bmatrix}=\begin{bmatrix}3 &-1 & -1 \\ 1 & 1 & -1 \\ 1 &-1 & 1 \end{bmatrix} \begin{bmatrix}x \\ y \\ z \end{bmatrix}"

The coefficient matrix has eigenvalues "\lambda" such that

"\begin{vmatrix} 
 3-\lambda & -1 & -1 \\
 1 & 1-\lambda & -1 \\
 1 & -1 & 1-\lambda \\
 \end{vmatrix} 
=-(\lambda-2)^2(\lambda-1)=0 \implies 2_{(2)}, \lambda=1"

(where the subscript denotes multiplicity of the eigenvalue).

"\lambda =2" :

"\begin{bmatrix}1 &-1 & -1 \\ 1 & -1 & -1 \\ 1 &-1 & -1 \end{bmatrix} \vec n_1=\vec 0\\
\implies n_{1,1}-n_{1,2}-n_{1,3}=0 \implies n_{1,1}=n_{1,2}+n_{1,3}"

By picking "n_{1,1}=1" , we can then set "n_{1,2}=1" and "n_{1,3}=0" , and vice versa, to find two corresponding eigenvectors,

"\vec n_1=\begin{bmatrix}1 \\ 1 \\ 0 \end{bmatrix}, \vec n_2=\begin{bmatrix}1 \\ 0 \\ 1 \end{bmatrix}"

"\lambda=1" :

"\begin{bmatrix}2 &-1 & -1 \\ 1 & 0 & -1 \\ 1 &-1 & 0 \end{bmatrix} \vec n_3=\vec 0\\
\implies 2n_{3,1}-n_{3,2}-n_{3,3}=0 \implies 2n_{3,1}=n_{3,2}+n_{3,3}"

We obtain "\vec n_3" independent of "\vec n_1, \vec n_2" by picking "\vec n_{3,2}=\vec n_{3,3}=1", so that the third corresponding eigenvector is;

"\vec n_3=\begin{bmatrix}1 \\ 1 \\ 1 \end{bmatrix}"

Then the general solution to this system is

"\begin{bmatrix}x \\ y \\ z \end{bmatrix}=c_1e^{2t} \vec n_1+ c_2te^{2t} \vec n_2+ c_3e^{t} \vec n_3\\
\begin{cases}
 x=c_1e^{2t} + c_2te^{2t} + c_3e^{t}\\ 
 y=c_1e^{2t} + c_3e^{t}\\ 
 z=c_2te^{2t} + c_3e^{t}\\ 
\end{cases}"